OpenStudy (dinorap1):

2. Solve 4x=28 by finding square roots. I don't know what this question wants? If it wants me to solve for x, then obviously x=7. But it's asking for square roots so...?

OpenStudy (anonymous):

that's super confusing. are you sure it's not x^2 or something??

OpenStudy (dinorap1):

No it's definitely 4x=28. I'm not sure what it wants me to do... I've been working on quadratic factoring lately, so maybe it has something to do with that?

OpenStudy (anonymous):

hmmm. so quadratic factoring. you could subtract 28 from both sides to get 4x-28=0, factor out 4 to get 4(x-7)=0 but that's still not using square roots, that's just a more roundabout way of doing it

OpenStudy (dinorap1):

Yeah I have no idea what it has to do with square roots. :/ 28 is not a perfect square, and x isn't squared either, so I have no clue.

OpenStudy (dinorap1):

Can you help me with this one instead @bigsby123 ? 3. Find the x-intercepts of y = −3x2 + 4x + 4 using the Quadratic Formula.

OpenStudy (anonymous):

sure. so do you know the quadratic formula?

OpenStudy (dinorap1):

Maybe you can refresh my memory on it. I'm not sure what it's asking for. What are x-intercepts?

OpenStudy (anonymous):

ok, so the quadratic formula is -b +/- sqrt ((b^2-4ac)/2a), where a=the coefficient next to x^2, b =coefficient next to x, and c=the number

OpenStudy (anonymous):

and the point of the quadratic formula is to find the x intercepts. the quadratic formula is set equal to x

OpenStudy (dinorap1):

Oh right. I remember that formula now. So I plug the numbers from #3 into that formula and x is my answer?

OpenStudy (anonymous):

correct. so tell me what you get as the x intercepts when you are finished. :)

OpenStudy (dinorap1):

Hahahahaha I just figured out what was wrong with #2 and #3! The exponent is there, but it's really tiny and transparent. So apparently number 2 is 4x^2=28.

OpenStudy (anonymous):

ok, that makes a lot more sense!

OpenStudy (dinorap1):

I will solve number 3 first and then go back to number 2.

OpenStudy (anonymous):

also, I had a typo: the quadratic formula is only square rooted over the numerator of the fraction, not the entire thing. sorry

OpenStudy (dinorap1):

@bigsby Okay here's what I got so far: x=-4(+-)8/6 Hopefully that makes sense. I had the correct formula written in my book already so don't worry. How do I simplify from this point?

OpenStudy (anonymous):

check your denominator again. remember because it is 2a, and a is -3, you're going to get a negative number

OpenStudy (anonymous):

other than that, you're correct!

OpenStudy (dinorap1):

Good thing you caught that! I had that written down but I forgot to write it on here. So is this what I would put as my final answer? Or could I simplify further?

OpenStudy (dinorap1):

I put it as: x=2/3 or x=-2/3

OpenStudy (dinorap1):

Nevermind it's x=-2/3 or x=2

OpenStudy (anonymous):

You are going to end up with two answers in the form of x=__ and ___. So if you choose the + path first, you would go through this process: -4+8/-6=-2/3. However, for the - path, it will be -4- 8/-6=2

OpenStudy (anonymous):

nice job!

OpenStudy (dinorap1):

Thanks! Now I just have to go back to number 2. 2. 4x^2=28

OpenStudy (anonymous):

Ok, so do you know where to start?

OpenStudy (dinorap1):

I think so. You subtract 28 over to the left side to get: 4x^2-28=0

OpenStudy (anonymous):

sure, that's one way to do it! so now what do you do?

OpenStudy (dinorap1):

I'm not sure. Can you help me remember what to do from here? I always have trouble factoring..

OpenStudy (anonymous):

I would factor out a 4 to get 4(x^2-7)=0. Do you know what to do next?

OpenStudy (dinorap1):

No I don't know what to do. :(

OpenStudy (anonymous):

ok, divide 4 by both sides. What do you have now?

OpenStudy (anonymous):

*on

OpenStudy (dinorap1):

I don't get when you're supposed to use square roots on this if 28 isn't a perfect square...

OpenStudy (dinorap1):

@bigsby123 Okay I'll try what you said. You say to divide 4 by both sides? What do you mean by that?

OpenStudy (dinorap1):

@bigsby123 Hello?

OpenStudy (anonymous):

so when you divide by 4 on both sides you're going to get x^2-7=0. Add 7 to both sides to get x^2=7. Then what do you do?

OpenStudy (anonymous):

en you divide by 4 on both sides you're going to get x^2-7=0. Add 7 to both sides to get x^2=7. Then what do you do?

OpenStudy (dinorap1):

Are you solving for x?

OpenStudy (anonymous):

yes, so this is when the square root comes in. square root both sides to get x= +/- sqrt 7

OpenStudy (anonymous):

got it?

OpenStudy (dinorap1):

Oh okay, so then the answer is sqrt 7. I thought you were supposed to factor for this question, but apparently not. THanks for the help!

OpenStudy (anonymous):

yup!

OpenStudy (dinorap1):

How do you do this one? 10. Use GCF to factor 15x^2+6x